OpenStudy (anonymous):

calculate the range for the following data set: 16,35,25,29,26,30,42,18

OpenStudy (amistre64):

1.68 2.596 3.50 4.2 is a histogram...

OpenStudy (amistre64):

let me ceck, but the range is from 16 to 42 so im thinking: 42-10 = 32 -6 = 26

OpenStudy (amistre64):

yep; range is the measure of variation that is the difference between the highest and lowest scores